# Data Augmentation for Supervised Graph Outlier Detection with Latent Diffusion Models

### v2

@@ Abstract
Graph outlier detection is a prominent task of research and application in the realm of graph neural networks. It identifies the outlier nodes that exhibit deviation from the majority in the graph. One of the fundamental challenges confronting supervised graph outlier detection algorithms is the prevalent issue of class imbalance, where the scarcity of outlier instances compared to normal instances often results in suboptimal performance. Conventional methods mitigate the imbalance by reweighting instances in the estimation of the loss function, assigning higher weights to outliers and lower weights to inliers. Nonetheless, these strategies are prone to overfitting and underfitting, respectively. Recently, generative models, especially diffusion models, have demonstrated their efficacy in synthesizing high-fidelity images. Despite their extraordinary generation quality, their potential in data augmentation for supervised graph outlier detection remains largely underexplored.
@@
To bridge this gap, we introduce GODM, a novel data augmentation for mitigating class imbalance in supervised Graph Outlier detection with latent Diffusion Models. Specifically, our proposed method consists of three key components: (1) Variantioanl Encoder maps the heterogeneous information inherent within the graph data into a unified latent space. (2) Graph Generator synthesizes graph data that are statistically similar to real outliers from latent space, and (3) Latent Diffusion Model learns the latent space distribution of real organic data by iterative denoising. Extensive experiments conducted on multiple datasets substantiate the effectiveness and efficiency of GODM. The case study further demonstrated the generation quality of our synthetic data. To foster accessibility and reproducibility, we encapsulate GODM into a plug-and-play package and release it at the Python Package Index (PyPI): https://pypi.org/project/godm/.
Index Terms:
Graph Outlier Detection, Data Augmentation, Diffusion Models, Class Imbalance, Anomaly Detection
I Introduction
Graphs are foundational data structures that capture complex relationships among diverse entities. Recently, graph data mining has showcased effectiveness and flexibility across various tasks, including community classification among academic citation networks [1], recommender systems in social networks [2], knowledge graph completion [3], and so on and forth. Within this broad spectrum of applications, graph outlier detection has emerged as a popular and important area of research and practice [4, 5, 6]. Graph outlier detection focuses on detecting outliers within graph-structured data that significantly deviate from standard patterns, which has proven valuable in domains such as financial fraud detection in bank systems [7], fake news detection on social media [8], spam detection on review platform [9], and anti-money laundering in transaction networks [10].
@@ I Introduction
However, despite the advancements in graph outlier detection techniques, similar to supervised outlier detection on other data modalities, graph outlier detection suffers a fundamental challenge known as class imbalance, where the number of positive instances (outliers) is significantly lower than the number of negative instances (inliers). For example, the ratio of positive to negative is only 1:85 in the DGraph dataset [7], reflecting the real-world extreme ratio in financial fraud detection scenario. This class imbalance problem poses challenges in the training of outlier detectors and often results in suboptimal performance. Specifically, as the negative instances dominate the training data, the loss function can be biased towards the majority of the negative class and, hence, exhibit poor generalization capability in identifying true outliers. In the worst case, a detector could classify every instance as negative, displaying seemingly low loss value and high accuracy score, yet basically lacking any knowledge.
@@
Figure 1: An toy example of data augmentation for class imbalance in graph outlier detection. Synthetic outliers help balance the class distribution in the training data, thereby improving the performance of graph outlier detection.
@@
Common practices mitigate this imbalance by upsampling or downsampling [11]. Upsampling augments the minority positive class by simply replicating outliers, whereas downsampling reduces the size of the majority negative class by randomly dropping normal instances. However, while beneficial in certain contexts, these methods often introduce their own challenges, such as the risk of overfitting outliers in the case of upsampling or the loss of valuable training data through downsampling. Another common approach to alleviate class imbalance is through instance reweighting in the loss function estimation. It allocates greater weights to positive instances and assigns lesser weights to negative instances. This approach is mathematically equivalent to upsampling and downsampling, thereby having a similar tendency to overfit outliers or underfitting inliers. These challenges underscore the need for more sophisticated data augmentation methods that can generate some synthetic data to handle the class imbalance in graph outlier detection, as Fig. 1 shows.
@@
In recent years, generative models, particularly diffusion models, have achieved significant advancements in synthesizing high-fidelity image data [12]. Diffusion models capture intricate data distributions and generate high-quality samples by gradually denoising the samples from a simple prior distribution (e.g., Gaussian distribution). On graph data, although some works have explored the possibility of diffusion on molecular graphs [13, 14], few studies have been conducted to apply diffusion models to tackle the class imbalance problem in the task of graph outlier detection. However, existing diffusion models can hardly be directly adapted to large-scale graph outlier detection applications due to the following non-trivial challenges: (1) Heterogeneity: Unlike relatively simple molecular graphs, the information contained in graphs for outlier detection can be highly heterogeneous, including high dimensional features and more than one type of edge, even temporal information. Existing diffusion models are primarily designed for monotypic information, exemplified by the RGB channels in image data. (2) Efficiency: In the application of graph outlier detection, e.g., financial fraud detection, the graphs are typically much larger than molecular graphs, up to millions, even billions scale. The prohibitive computational cost of diffusion models hinders their direct application in outlier detection on large graphs. (3) Condition: Different from the unconditional generation on molecular graphs, conditional generation is required to mitigate the problem of class imbalance. We only want to generate outliers rather than normal nodes in synthetic graphs.
@@
To bridge this gap, we make the first attempt to leverage diffusion models in graph outlier detection and propose a data augmentation for mitigating class imbalance in supervised Graph Outlier detection with latent Diffusion Models (GODM). Our main idea is to generate outliers in graph space while conducting diffusion in latent space. To address heterogeneity, we propose Variantioanl Encoder to map the heterogeneous information inherent within the graph data into a unified latent space. In addition, Graph Generator synthesizes different types of information back to graph space from the latent embedding. To alleviate the efficiency problem, instead of direct diffusion in graph space, we only conduct diffusion in the latent space crafted by the variational encoder. Furthermore, we use negative sampling and graph clustering to reduce the computational cost. For diffusion models, we also adopt EDM [15] instead of commonly used DDPM [16] to facilitate the generation efficiency of Latent Diffusion Model. For the condition, we not only give a class label to the variational encoder to form node embedding with class information but also conduct conditional generation on both Latent Diffusion Model and Graph Generator. Finally, our heterogeneous, efficient, and conditional GODM can generate graphs with outliers that are integrated with the original real graph for the training of the downstream graph outlier detector. Importantly, GODM is model agnostic, providing researchers and practitioners with the flexibility to integrate it across various graph outlier detectors.
Our contributions in this paper are mainly as follows:
  • We make the first attempt to leverage diffusion models in graph outlier detection and propose to adopt diffusion models in the latent space and bridge the graph space and the latent space with a variational encoder and graph generator for heterogeneity.
  • We further design outlier-only generation conditioned on outlier labels and improve the efficiency of GODM by negative sampling, graph clustering, and EDM.
  • We conduct extensive experiments to demonstrate the effectiveness and efficiency of GODM. To foster accessibility, we make our code a plug-and-play package, which is publicly available at: https://pypi.org/project/godm/.
The remainder of this paper is organized as follows. Section II formally defines the notation and the problem we studied in this paper. Section III provides an overview of GODM. Section IV gives detailed experimental results and analyses. Section V briefly describes related works. We summarize the paper and discuss future work in Section VI.
II Preliminary
Figure 2: The architecture of proposed data augmentation method GODM. For a given graph $\mathcal{G}$, GODM first partitions the large graph by graph clustering algorithms and conducts negative sampling on edge, reducing the computational cost. Then, a GNN-based Variational Encoder maps each node in subgraphs into the latent space embedding $\mathbf{Z}^{0}$. A forward diffusion process $\mathbf{Z}^{0}\rightarrow\mathbf{Z}^{S}$ is applied in the latent space. Reverse denoising process $\hat{\mathbf{Z}}^{S}\rightarrow\hat{\mathbf{Z}}^{0}$ starts from the prior distribution $p(\hat{\mathbf{Z}}^{S})$ and generates samples $\hat{\mathbf{Z}}^{0}$ in latent space with denoising MLP conditioned on outlier labels $\hat{\mathbf{y}}$. Graph Generator also conditionally generate node features $\hat{\mathbf{X}}$, graph structures $\hat{\mathcal{E}}$, edge timestamps $\hat{\mathbf{t}}$, and edge types $\hat{\mathbf{p}}$ from latent embedding $\hat{\mathbf{Z}}^{0}$. By ensembling all necessary information, we obtain the synthetic graph $\hat{\mathcal{G}}$. Along with the real graph ${\mathcal{G}}$, $\hat{\mathcal{G}}$ will be used for the training of the downstream graph outlier detector $D$.
@@ II Preliminary
II-A Notation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X},\mathbf{y},\mathbf{t},\mathbf{p})$ denotes a graph with $n$ nodes, where $\mathcal{V}=\{v_{i}\}_{i=1}^{n}$ is the set of nodes, and =$\mathcal{E}=\{e_{ij}\}$ represents the set of edges. Here, $e_{ij}=(v_{i},v_{j})$ denotes an edge between node $v_{i}$ and node $v_{j}$. The matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ contains $d$-dimensional feature vectors $\mathbf{x}_{i}$ for each node $v_{i}$.  $\mathbf{y}\in\{0,1\}^{n}$ represents the vector of the node label $y_{i}$ for each node $v_{i}$, where 00 denotes an inlier, while $1$ denotes an outlier. $\mathbf{t}=\{t_{ij}\}\in\mathbb{N}$ and $\mathbf{p}=\{p_{ij}\}\in\{1,\ldots,P\}$ are the optional non-negative integer edge timestamp vector and edge type vector, respectively, where $P$ is the number of edge types.
II-B Problem Formulation
@@
Given a graph $\mathcal{G}$, graph outlier detection can be regarded as a binary classification task that learns a detector $D:v_{i}\rightarrow\{0,1\}$ that classifies every node in $\mathcal{G}$ to an inlier (0) or an outlier (1).
@@
In outlier detection, the number of inliers is far more than the number of outliers, i.e., |$|\{v_{i}\mid y_{i}=0\}|\gg|\{v_{i}\mid y_{i}=1\}|$. Therefore, we aim to learn a parameterized data augmentation model, which can generate realistic and diverse synthetic graph $\hat{\mathcal{G}}=(\hat{\mathcal{V}},\hat{\mathcal{E}},\hat{\mathbf{X}},\hat{\mathbf{y}},\hat{\mathbf{t}},\hat{\mathbf{p}})$, where $\hat{y_{i}}=1,\forall v_{i}\in\hat{\mathcal{V}}$. To alleviate the class imbalance in the training of the graph outlier detector $D$, we integrate the original graph $\mathcal{G}$ and the synthetic graph $\hat{\mathcal{G}}$ for the training of detector $D$.
III Methodology
@@
In this section, we elaborate on the proposed data augmentation method GODM in detail. Fig. 2 shows the architecture of GODM. It starts by partitioning the input graph $\mathcal{G}$ with graph clustering algorithms for reducing memory consumption. Subsequently, each partitioned subgraph is encoded into a latent space representation denoted by $\mathbf{Z}^{0}$ through Variational Encoder based on graph neural networks (GNNs), followed by a forward diffusion $\mathbf{Z}^{0}\rightarrow\mathbf{Z}^{S}$ in the latent space, where $S$ is the maximum step in the diffusion model. The reverse denoising process $\hat{\mathbf{Z}}^{S}\rightarrow\hat{\mathbf{Z}}^{0}$ iteratively estimates the noise and generates the latent embedding $\hat{\mathbf{Z}}^{0}$ from a predefined prior distribution $p(\hat{\mathbf{Z}}^{S})$ using a denoising function, which is a multi-layer perceptron (MLP), conditioned on class labels $\hat{\mathbf{y}}$. Graph Generator employs the latent representations $\hat{\mathbf{Z}}^{0}$ to conditionally reconstruct the node features and graph structures alongside other information available in the original real graph. By ensembling generated information, we are able to obtain $\hat{\mathcal{G}}$ that contains the nodes that are statistically similar to the real outliers. This synthetic graph $\hat{\mathcal{G}}$, in concert with the original graph $\mathcal{G}$, is leveraged to enhance the training efficacy of the downstream graph outlier detection task.
In Section III-A and Section III-B, we introduce the design details of the Variational Encoder and Graph Generator, respectively. Then, we describe Latent Diffusion Model in Section III-C. We summarize the training and inference process in Section III-D and Section III-E, respectively. Finally, we analyze the algorithmic complexity in Section III-F.
III-A Variational Encoder
@@
Graph data is inherently complex and heterogeneous, containing both node feature $\mathbf{X}$ and graph structure $\mathcal{E}$, sometimes also partial node class labels $\mathbf{y}$, temporal information $\mathbf{t}$, and different types of edges $\mathbf{p}$. However, current diffusion models are mainly designed for monotypic information (e.g., the magnitude of RGB channels in images). To bridge this gap, we adopt a GNN-based Variational Encoder $E:\mathcal{G}\rightarrow\mathbf{Z}$ to map different types of information to a unified latent space.
@@
Node Feature. To encode the node feature into the latent space, we take the feature of each node as the initial embedding for the encoder, i.e., $\mathbf{h}_{i}^{0}=\mathbf{x}_{i}$.
@@
Class Label. As we only generate outliers by conditional generation, it will be helpful to encode node labels $\mathbf{y}=\{y_{i}\}$. We add the class information to the initial node embedding:
$\mathbf{h}_{i}^{0}=\mathbf{x}_{i}+\mathbf{w}^{C}_{E}\cdot y_{i},$ (1)
where $\mathbf{w}^{C}_{E}\in\mathbb{R}^{d}$ is a linear transformation for class label.
@@
Graph Structure. Graph neural networks have emerged as a profound architecture for learning graph data, efficaciously learning the node representations by encoding the graph topology and node feature simultaneously with message passing. In this paper, we take GraphSAGE [17] as an example. For each layer of GraphSAGE, each node updates its embedding by aggregating the message within the neighborhood:
@@
where $\mathbf{h}_{i}^{l}$ is the node embedding for node $v_{i}$ in the $l-$th layer, $\mathbf{m}_{ij}^{l}$ is the edge-wise message of edge $e_{ij}$, and ACT is a non-linear activation function. $\mathbf{W}^{l}\in\mathbb{R}^{d^{l}\times 2d^{l-1}}$ is the linear transformation matrix, where $d^{l}$ and $d^{l-1}$ are the hidden dimension of $l$-th layer and $(l-1)$-th layer, respectively. While CAT represents concatenation, AGG denotes the aggregation function (e.g., mean). In the vanilla GraphSAGE, an edge-wise message is typically the embedding of source node $v_{j}$ from the last layer:
$\mathbf{m}_{ij}^{l}=\mathbf{h}_{j}^{l-1}.$ (3)
Edge Type. However, sometimes, we have different types of edges in graphs, known as heterogeneous graphs or heterogeneous information networks [18, 19]. As different edge types can encapsulate various semantics, encoding edge types can be important for the downstream task. Therefore, we add the type information to the edge-wise message:
$\mathbf{m}_{ij}^{l}=\mathbf{h}_{j}^{l-1}+\mathbf{W}^{P}\cdot\phi(p_{ij}),$ (4)
where $\mathbf{W}^{P}\in\mathbb{R}^{d^{l-1}\times P}$ is a linear transformation for edge type, and $\phi$ is the one-hot encoding function.
Edge Time. In addition, temporal information is critical in time series applications [20]. When edge time is available, we are also able to encode the timestamp to the edge-wise message with trigonometric temporal embedding [21]:
$\mathbf{m}_{ij}^{l}=\mathbf{h}_{j}^{l-1}+\mathbf{W}^{T}\cdot TE_{(t_{ij},\cdot)},$ (5)
where $\mathbf{W}^{T}\in\mathbb{R}^{d^{l-1}\times d^{l-1}}$ is the linear transformation for edge time, and $TE$ is defined as:
$TE_{(t_{ij},2k)}=sin(t_{ij}/10000^{2k/d^{l-1}}),$ (6)
$TE_{(t_{ij},2k+1)}=cos(t_{ij}/10000^{2k/d^{l-1}}),$
where $k$ is each dimension of the temporal embedding.
By stacking multiple layers, GNNs are capable of encoding both node features and neighborhood graph structures into independent and identically distributed (i.i.d.) node embedding $\mathbf{Z}$, which can be further leveraged by Latent Diffusion Model. To ease the generation for Latent Diffusion Model, we use Variational Encoder, which outputs two matrixes:
$\bm{\mu}=\text{GNN}_{\bm{\mu}}(\text{GNN}_{\text{shared}}(\mathcal{G})),$ (7)
$\text{log}\bm{\sigma}=\text{GNN}_{\bm{\sigma}}(\text{GNN}_{\text{shared}}(\mathcal{G})),$
@@
$\mathbf{Z}=\bm{\mu}+\bm{\sigma}\cdot\bm{\varepsilon},\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I}),$ (8)

### v3

# Data Augmentation for Supervised Graph Outlier Detection via Latent Diffusion Models

@@ Abstract
A fundamental challenge confronting supervised graph outlier detection algorithms is the prevalent problem of class imbalance, where the scarcity of outlier instances compared to normal instances often results in suboptimal performance. Recently, generative models, especially diffusion models, have demonstrated their efficacy in synthesizing high-fidelity images. Despite their extraordinary generation quality, their potential in data augmentation for supervised graph outlier detection remains largely underexplored. To bridge this gap, we introduce GODM, a novel data augmentation for mitigating class imbalance in supervised Graph Outlier detection via latent Diffusion Models. Extensive experiments conducted on multiple datasets substantiate the effectiveness and efficiency of GODM. The case study further demonstrated the generation quality of our synthetic data. To foster accessibility and reproducibility, we encapsulate GODM into a plug-and-play package and release it at PyPI: https://pypi.org/project/godm/.
1 Introduction
@@
Graph outlier detection has emerged as a popular and important area of research and practice in graph machine learning [1, 2, 3]. Graph outlier detection focuses on detecting outliers within graph-structured data that significantly deviate from standard patterns. It has been proven valuable in domains such as fraud detection [4], fake news detection [5], spam detection [6], anti-money laundering [7], etc. Despite the advancements in graph outlier detection techniques, similar to supervised outlier detection on other data modalities, graph outlier detection suffers a fundamental challenge known as class imbalance. This challenge manifests as a significantly lower number of positive instances (outliers) compared to negative instances (inliers). For example, the ratio of positive to negative is only 1:85 on DGraph dataset [4], reflecting the extreme ratio in real-world financial fraud detection scenarios. This class imbalance problem poses challenges in the training of graph outlier detectors and often results in suboptimal performance. Specifically, as the negative instances dominate the training data, the loss function can be biased towards the majority of the negative class and, hence, exhibit poor generalization capability in identifying true outliers.
@@
Figure 1: An toy example of data augmentation for class imbalance in graph outlier detection.
@@
Common practices mitigate this imbalance by upsampling or downsampling [8]. Upsampling augments the minority positive class by replicating outliers, whereas downsampling reduces the size of the majority negative class by randomly dropping normal instances. However, these methods often have their own risks, such as overfitting outliers in upsampling or the loss of valuable training data through downsampling. Another common approach to alleviate class imbalance is through instance reweighting in the loss function, assigning greater weights to outliers and less weights to inliers. This approach is mathematically equivalent to upsampling and downsampling, thereby having a similar problem. For a more comprehensive summary of previous methods, please refer to Appendix A. These challenges underscore the need for more sophisticated data augmentation methods that can generate some synthetic data to balance the class distribution in the training data, thereby improving the performance of graph outlier detection, as Figure 1 shows.
@@
In recent years, generative models, particularly diffusion models, have achieved significant advancements in synthesizing high-fidelity image data [9]. Diffusion models capture intricate data distributions and generate high-quality samples by gradually denoising the samples from a simple prior distribution (e.g., Gaussian distribution). Diffusion models provide more generative power compared to variational autoencoders and offer more stable training than generative adversarial networks. On graph data, although some works have explored the possibility of diffusion on molecular graphs [10, 11], few studies have been conducted to apply diffusion models to tackle the class imbalance problem in the task of graph outlier detection. However, existing diffusion models can hardly be directly adapted to large-scale graph outlier detection applications due to the following non-trivial challenges: (1) Heterogeneity: Unlike relatively simple molecular graphs, the information contained in graphs for outlier detection can be highly heterogeneous, including high dimensional features and more than one type of edge, even temporal information. Existing diffusion models are primarily designed for monotypic information, exemplified by the RGB channels in image data. (2) Efficiency: In the application of graph outlier detection, e.g., financial fraud detection, the graphs are typically much larger than molecular graphs, up to millions, even billions scale. The prohibitive computational cost of diffusion models hinders their direct application in outlier detection on large graphs. (3) Condition: As we only want to generate outliers rather than normal nodes in synthetic graphs, conditional generation is required to mitigate the problem of class imbalance.
@@
To bridge this gap, we apply diffusion models in graph outlier detection and propose a data augmentation for mitigating class imbalance in supervised Graph Outlier detection via latent Diffusion Models (GODM). Our main idea is to generate outliers in graph space while conducting diffusion in latent space. To address heterogeneity, we propose Variational Encoder to map the heterogeneous information inherent within the graph data into a unified latent space. In addition, Graph Generator synthesizes different types of information back to graph space from the latent embedding. To alleviate the efficiency problem, instead of direct diffusion in graph space, we only conduct diffusion in the latent space crafted by the variational encoder. Furthermore, we use negative sampling and graph clustering to reduce the computational cost. For diffusion models, we also adopt EDM [12] instead of commonly used DDPM [13] to facilitate the generation efficiency of Latent Diffusion Model. For the condition, we not only give a class label to the variational encoder to form node embedding with class information but also conduct conditional generation on both Latent Diffusion Model and Graph Generator. Finally, our heterogeneous, efficient, and conditional GODM can generate graphs with outliers that are integrated with the original real graph for the training of the downstream graph outlier detector. Importantly, GODM is model agnostic, providing researchers and practitioners with the flexibility to integrate it across various graph outlier detectors. To foster accessibility, we make our code a plug-and-play package, which can be easily adopted on PyG Data object.
2 Preliminary
@@
2.1 Notation
Let $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X},\mathbf{y},\mathbf{t},\mathbf{p})$ denotes a graph with $n$ nodes, where $\mathcal{V}=\{v_{i}\}_{i=1}^{n}$ is the set of nodes, and =$\mathcal{E}=\{e_{ij}\}$ represents the set of edges. Here, $e_{ij}=(v_{i},v_{j})$ denotes an edge between node $v_{i}$ and node $v_{j}$. The matrix $\mathbf{X}\in\mathbb{R}^{n\times d}$ contains $d$-dimensional feature vectors $\mathbf{x}_{i}$ for each node $v_{i}$.  $\mathbf{y}\in\{0,1\}^{n}$ represents the vector of the node label $y_{i}$ for each node $v_{i}$, where 00 denotes an inlier, while $1$ denotes an outlier. $\mathbf{t}=\{t_{ij}\}\in\mathbb{N}$ and $\mathbf{p}=\{p_{ij}\}\in\{1,\ldots,P\}$ are the optional non-negative integer edge timestamp and edge type, respectively, where $P$ is the number of edge types.
2.2 Problem Formulation
@@
In this paper, we focus on node level outlier detection. Given a graph $\mathcal{G}$, graph outlier detection can be regarded as a binary classification task that learns a detector $D:v_{i}\rightarrow\{0,1\}$ that classifies every node in $\mathcal{G}$ to an inlier (0) or an outlier (1).
@@
In outlier detection, the number of inliers is far more than the number of outliers, i.e., |$|\{v_{i}\mid y_{i}=0\}|\gg|\{v_{i}\mid y_{i}=1\}|$. We aim to learn a parameterized data augmentation model to generate realistic and diverse synthetic graph $\hat{\mathcal{G}}=(\hat{\mathcal{V}},\hat{\mathcal{E}},\hat{\mathbf{X}},\hat{\mathbf{y}},\hat{\mathbf{t}},\hat{\mathbf{p}})$, where $\hat{y_{i}}=1,\forall v_{i}\in\hat{\mathcal{V}}$. The synthetic graph $\hat{\mathcal{G}}$ is integrated with the original graph $\mathcal{G}$ to alleviate the class imbalance in the training of the graph outlier detector $D$.
3 Methodology
Figure 2: The architecture of proposed data augmentation method GODM.
In this section, we elaborate on the proposed data augmentation method GODM.
3.1 Overview
@@
Figure 2 shows the architecture of GODM. It starts by partitioning the input graph $\mathcal{G}$ with graph clustering algorithms and conducting negative sampling on edges to reduce memory consumption. Subsequently, each partitioned subgraph is encoded into a latent space representation denoted by $\mathbf{Z}^{0}$ through Variational Encoder based on graph neural networks (GNNs), followed by a forward diffusion $\mathbf{Z}^{0}\rightarrow\mathbf{Z}^{S}$ in the latent space, where $S$ is the maximum step in the diffusion model. The reverse denoising process $\hat{\mathbf{Z}}^{S}\rightarrow\hat{\mathbf{Z}}^{0}$ iteratively estimates the noise and generates the latent embedding $\hat{\mathbf{Z}}^{0}$ from a predefined prior distribution $p(\hat{\mathbf{Z}}^{S})$ using a denoising function, which is a multi-layer perceptron (MLP), conditioned on class labels $\hat{\mathbf{y}}$. Graph Generator employs the latent representations $\hat{\mathbf{Z}}^{0}$ to conditionally reconstruct the node features and graph structures alongside other information available in the original real graph. By ensembling generated information, we are able to obtain $\hat{\mathcal{G}}$ that contains the nodes that are statistically similar to the real outliers. This synthetic graph $\hat{\mathcal{G}}$, in concert with the original graph $\mathcal{G}$, is leveraged to the training of the downstream graph outlier detector $D$.
3.2 Variational Encoder
@@
Graph data is inherently complex and heterogeneous, containing both node feature $\mathbf{X}$ and graph structure $\mathcal{E}$, sometimes also partial node labels $\mathbf{y}$, temporal information $\mathbf{t}$, and edge types $\mathbf{p}$. However, current diffusion models are mainly designed for monotypic information (e.g., the magnitude of RGB channels in images). To bridge this gap, we adopt a GNN-based Variational Encoder $E:\mathcal{G}\rightarrow\mathbf{Z}$ to map different types of information in the graph space into a unified latent space.
@@
Node Feature. To encode the node feature into the latent space, we take the feature of each node as the initial embedding for the input of the encoder, i.e., $\mathbf{h}_{i}^{0}=\mathbf{x}_{i}$.
@@
Class Label. As we only generate outliers by conditional generation, we encode node labels $\mathbf{y}=\{y_{i}\}$ into the initial node embedding by: $\mathbf{h}_{i}^{0}=\mathbf{x}_{i}+\mathbf{w}^{C}_{E}\cdot y_{i}$, where $\mathbf{w}^{C}_{E}\in\mathbb{R}^{d}$ is a linear transformation.
@@
Graph Structure. Graph neural networks (GNNs) have emerged as a profound architecture for learning graph data. GNNs efficaciously learn the node representations by encoding the graph topology and node feature simultaneously with message passing. In this paper, we take GraphSAGE [14] as an example. For each layer of GraphSAGE, each node updates its embedding by aggregating the message within the neighborhood:
@@
where $\mathbf{h}_{i}^{l}$ is the node embedding for node $v_{i}$ in the $l-$th layer, $\mathbf{m}_{ij}^{l}$ is the edge-wise message of edge $e_{ij}$, and ACT is a non-linear activation function. $\mathbf{W}^{l}\in\mathbb{R}^{d^{l}\times 2d^{l-1}}$ is the linear transformation, where $d^{l}$ and $d^{l-1}$ are the hidden dimension of $l$-th layer and $(l-1)$-th layer, respectively. While CAT represents concatenation, AGG denotes the aggregation function (e.g., mean). In vanilla GraphSAGE, an edge-wise message is the embedding of source node $v_{j}$ from the last layer $\mathbf{m}_{ij}^{l}=\mathbf{h}_{j}^{l-1}$.
Edge Type. However, sometimes, we have different types of edges in graphs, known as heterogeneous graphs or heterogeneous information networks [15, 16]. As different edge types can encapsulate various semantics, encoding edge types can be important for the downstream task. Therefore, we add the edge type information to the edge-wise message $\mathbf{m}_{ij}^{l}=\mathbf{h}_{j}^{l-1}+\mathbf{W}^{P}\cdot\phi(p_{ij})$, where $\mathbf{W}^{P}\in\mathbb{R}^{d^{l-1}\times P}$ is a linear transformation for edge type, and $\phi$ is the one-hot encoding.
Edge Time. In addition, temporal information is critical in time series applications [17]. When edge time is available, we are also able to encode the timestamp to the edge-wise message with temporal embedding: $\mathbf{m}_{ij}^{l}=\mathbf{h}_{j}^{l-1}+\mathbf{W}^{T}\cdot TE{(t_{ij},\cdot)}$, where $\mathbf{W}^{T}\in\mathbb{R}^{d^{l-1}\times d^{l-1}}$ is the linear transformation for edge time, and $TE$ is the trigonometric temporal embedding [18].
By stacking multiple layers, GNNs are capable of encoding both node features and neighborhood graph structures into independent and identically distributed (i.i.d.) node embedding $\mathbf{Z}$, which can be further leveraged by Latent Diffusion Model. To ease the generation for Latent Diffusion Model, we use Variational Encoder, which outputs two different matrixes:
$\bm{\mu}=\text{GNN}_{\bm{\mu}}(\text{GNN}_{\text{shared}}(\mathcal{G})),\quad\text{log}\bm{\sigma}=\text{GNN}_{\bm{\sigma}}(\text{GNN}_{\text{shared}}(\mathcal{G})),$ (2)
@@
$\mathbf{Z}=\bm{\mu}+\bm{\sigma}\cdot\bm{\varepsilon},\quad\bm{\varepsilon}\sim\mathcal{N}(\bm{0},\bm{I}),$ (3)